# Social Norms in Cinema: A Cross-Cultural Analysis of Shame, Pride and Prejudice

### v3

@@ A.7 Clustering Results
Inappropriate behavior not living up to someone’s expectations, being disloyal, Being rude and ungrateful
@@
Sexual Harrassment assaulting a girl, physical assault on women, sexual assault
@@
Alcholism excessive drinking, inappropriate behavior due to alcohol, drinking and irresponsible behavior
@@
Table A11: Themes and top-3 examples for clusters obtained after agglomerative clustering of shame-related dialogues from Bollywood and Hollywood.
@@
Bravery & Nation bravery and sacrifice, bravery and service to nation, bravery and selflessness
@@
Son’s success son’s success and progress, son’s hard work and achievement, son’s determination
@@
Marriage-related marrying with pride and respect, sister’s marriage, daughter’s marriage into a reputed family
Appearance & Assets his manhood, his wealth and beautiful wife, his knowledge and power
@@
Table A12: Themes and top-3 examples for clusters obtained after agglomerative clustering of pride-related dialogues from Bollywood and Hollywood.
@@
Table A13: Distribution of reasons (shame) across manually labeled clusters. A total of twenty-six clusters were generated with distance=5. Duplicates were removed for clustering. One cluster with generic reasons ( such as phrases "lack of shame") was removed and two clusters with similar reasons (related to accountability) were merged. Finally, 24 clusters were considered. The total is slightly more than numbers in Table A3 since a norm could be mapped to multiple clusters.
@@
Table A14: Distribution of reasons (pride) across manually labeled clusters. A total of fifteen clusters were generated with distance=5. Duplicates were removed for clustering. ’Achievement’ and ’Winning’ were merged to form ’Achievement’ and, ’Bravery’ and ’Nation’ were merged to form ’Bravery & Nation’. Finally, 13 clusters were considered for analysis.
(a) Shame-related norms in Bollywood
(b) Shame-related norms in Hollywood
Figure A1: Gender-wise differences in themes associated with normative expectations from shame-related discourse in Bollywood and Hollywood. $\Delta(Female,Male)$ is the difference between normalized dialogues attributed to females under a theme and to males under the same theme (as in eq. 1 but for dimension gender). * indicates insignificant difference i.e., $p>0.05$
(a) Shame-related norms in Bollywood
(b) Shame-related norms in Hollywood
Figure A2: Gender-wise differences in themes associated with normative expectations from pride-related discourse in Bollywood and Hollywood. $\Delta(Female,Male)$ is the difference between normalized dialogues attributed to females under a theme and to males under the same theme (as in eq. 1 but for dimension gender). * indicates insignificant difference i.e., $p>0.05$

### v4

@@ A.5 Clustering Results
Inappropriate social behavior not living up to someone’s expectations, being disloyal, Being rude and ungrateful
@@
Sexual Harassment assaulting a girl, physical assault on women, sexual assault
@@
Alcoholism excessive drinking, inappropriate behavior due to alcohol, drinking and irresponsible behavior
@@
Lack of shame perceived shamelessness, being perceived as shameless, feeling embarrassed and ashamed
Table A11: Themes ($n=26$) and top-3 examples for clusters obtained after agglomerative clustering of shame-related dialogues from Bollywood and Hollywood. The cluster Lack of shame was removed since the reason behind shame was not evident. Clusters Lack of accountability and Inappropriate social behavior were merged together (and named ‘‘Lack of accountability’’) due to overlapping normative expectations.
@@
Bravery bravery and sacrifice, bravery and service to nation, bravery and selflessness
Nation saving the country’s pride, making country proud, Making country proud
@@
Son’s achievements son’s success and progress, son’s hard work and achievement, son’s determination
@@
Wedding marrying with pride and respect, sister’s marriage, daughter’s marriage into a reputed family
Physical Appearance & Assets his manhood, his wealth and beautiful wife, his knowledge and power
@@
Winning Sports achievement, working hard and achieving a spectacular result, winning something
@@
Table A12: Themes ($n=15$) and top-3 examples for clusters obtained after agglomerative clustering of pride-related dialogues from Bollywood and Hollywood.
@@
Table A13: Distribution of reasons (shame) across manually labeled clusters. A total of twenty-six clusters were generated with distance=5. Duplicates were removed for clustering. One cluster with generic reasons ( such as phrases "lack of shame") was removed and two clusters with similar reasons (related to accountability) were merged. Finally, 24 clusters were considered. The total is slightly more than numbers in Table 5 since a reason could be mapped to multiple clusters.
@@
Table A14: Distribution of reasons (pride) across manually labeled clusters. A total of fifteen clusters were generated with distance=5. The total is slightly more than the numbers in Table 5 since a reason could be mapped to multiple clusters
(a) Pride-related norms in Bollywood
(b) Pride-related norms in Hollywood
Figure A1: Gender-wise differences in themes associated with normative expectations from pride-related discourse in Bollywood and Hollywood. $\Delta(Female-Male)$ is the difference between normalized dialogues attributed to females under a theme and to males under the same theme (as in eq. 1 but for dimension gender). * indicates significant difference i.e., $p>0.05$
Theme Statistic p-value
Inappropriate Sexual behavior 5.0380 0.00001
Gender roles 3.8262 0.0006
Betrayal 3.2044 0.001
Illegal activities 2.8258 0.005
Disrespect 2.8284 0.005
Sexual Harassment 2.8358 0.005
Promiscuity 1.5553 0.1
Accusation 3.8411 0.0006
Stealing 1.6798 0.09
Parent-related 1.8961 0.06
Family norms 2.0813 0.04
Non-conformity 1.4724 0.1
Marriage-related 1.1817 0.2
Privacy-related 0.9518 0.3
Alcoholism -0.2920 0.8
Immodesty -0.6634 0.5
Disobedience -2.1438 0.03
Lying/Deception -2.6508 0.008
Cowardice -2.7746 0.006
Incompetence -3.5187 0.002
Social Etiquette -3.3461 0.002
Harm -4.5140 0.00009
Poverty -4.9028 0.00002
Lack of accountability -4.6312 0.00007
Table A15: Barnard Exact test with Yates correction for testing statistically significant occurrence of shame-related themes in Bollywood and Hollywood.
Theme Statistic p-value
Ethnolinguistic Identity 9.6185 0.0000
Son’s Achievements 5.8025 0.0000
Bravery 5.0368 0.0000
Family Honor 4.7046 0.0000
Nation 5.1183 0.0000
Wedding 5.3898 0.0000
Resilience 1.4959 0.1370
Justice 0.1441 0.8963
Family Roles -0.2560 0.8024
Physical Appearance -0.6420 0.5223
Winning -2.8913 0.0038
Achievement -3.5425 0.0004
Self-identity -4.9482 0.0000
Doing the "right" thing -10.4936 0.0000
Duty -8.5488 0.0000
Table A16: Barnard Exact test with Yates correction for testing statistically significant occurrence of pride-related themes in Bollywood and Hollywood.